# V2Xum-LLM: Cross-Modal Video Summarization with Temporal Prompt Instruction Tuning

### v1

@@ 7 Conclusion
In this study, we address the deficiencies in current video summarization datasets including the insufficient number of training examples and the insufficient evaluation of video summarization by building a new large-scale cross-model video summarization dataset Instruct-V2Xum and designing the improved video summarization evaluation metrics. We also propose a novel temporal prompt instruction tuning method – V2Xum-LLM that unifies the generation of different modalities of video summaries into the text decoder of VLMs. This method does not rely on task-specific heads to perform video summarization tasks and enables the models to accept the interleaved long video and language input sequences. In addition, our proposed method also supports modality controllable summary generation through language instructions. The experimental results show the effectiveness of our proposed method.
Appendix A Importance Score Prediction with LLM Logits
@@
Appendix B Quality Analysis
We analyze the quality of GPT-Synthesised data from the perspective of grammar fluency and commonsense plausibility. For the grammar-fluency evaluation, we leverage a grammar-check model [46] that assigns high scores to correct texts grammatically. For the commonsense plausibility evaluation, we utilize Vera [40], a plausibility estimation model, to identify the nonsensical bias. The results are presented in Figure 4 and Figure 5. We can conclude that our dataset with the refinement process achieves a satisfactory score.
@@ Appendix B Quality Analysis
Appendix C Example Data from Instruct-V2Xum

### v2

@@ Conclusion
In this study, we address the deficiencies in current video summarization datasets including the insufficient number of training examples and the insufficient evaluation of video summarization by building a new large-scale cross-model video summarization dataset Instruct-V2Xum and designing the improved video summarization evaluation metrics. We also propose V2Xum-LLM, a novel temporal prompt instruction tuning method that unifies the generation of various video summary modalities within the text decoder of VLMs, eliminating the need for task-specific heads. This approach supports interleaved long video and language input sequences and allows modality-controllable summary generation through language instructions. Experimental results demonstrate the effectiveness of our method.
Appendix A Appendix
Importance Score Prediction with LLM Logits
@@ Appendix A Appendix
Quality Analysis
We analyze the quality of GPT-Synthesised data from the perspective of grammar fluency and commonsense plausibility. For the grammar-fluency evaluation, we leverage a grammar-check model (Morris et al. 2020) that assigns high scores to correct texts grammatically. For the commonsense plausibility evaluation, we utilize Vera (Liu et al. 2023b), a plausibility estimation model, to identify the nonsensical bias. The results are presented in Figure 3 and Figure 4. We can conclude that our dataset with the refinement process achieves a satisfactory score.
@@
Example Data from Instruct-V2Xum
@@
V2Xum-LLaMA Prediction Visualization
In this section, we visualize some results predicted by V2Xum-LLaMA in Figure 7.
Figure 7: Some results predicted by V2Xum-LLaMA.